OpenStudy (anonymous):

how do you solve the initial value problem of 4y(3)+2y=0;y(0)=2y(0)=5y(0)=6

OpenStudy (anonymous):

i got the wrong answer :(

OpenStudy (anonymous):

Can you retype the question? It didn't come out right? What are those 0s? Are they exponents or numbers?

OpenStudy (anonymous):

they are numbers

OpenStudy (anonymous):

But they are 0s?

OpenStudy (lalaly):

What course are studying?

OpenStudy (anonymous):

Here's is the problem again: 4y'''+2y''=0 y(0)=2 y'(0)=5 y''(0)=6

OpenStudy (anonymous):

yes they are zeros

OpenStudy (anonymous):

calc?

OpenStudy (lalaly):

differential eqns:D

OpenStudy (anonymous):

and im taking differential equations

OpenStudy (anonymous):

yes lol

OpenStudy (anonymous):

the 3 in the problem represents the third derivative im sorry the original post looks weird

OpenStudy (anonymous):

its supposed to be 4y'''+2y''=0 as in to the third derivative and the second derivative

OpenStudy (lalaly):

\[4r^3+2r^2=0\]\[2r^2(r+2)=0\]\[r=0,0,-1/2\]so \[y=c_1e^{-0.5x}+c_2x+c3\]

OpenStudy (lalaly):

\[y'=-0.5c_1e^{-0.5x}+c_2\]

OpenStudy (lalaly):

\[y''=0.25c_1e^{-0.5x}\]

OpenStudy (anonymous):

yes i got that but i think i did something wrong after that

OpenStudy (lalaly):

just plug in the values of x and y, nothing'd go wrong

OpenStudy (anonymous):

i did but the answer is wrong! i did y(0)=2=c_1 +c_3

OpenStudy (anonymous):

y'(0)=5=c2-(1/2)c3e^(-1/2x)

OpenStudy (lalaly):

no the e part cancels coz it equals 1 when x =0

OpenStudy (lalaly):

y'(0)=c2-1/2c3

OpenStudy (anonymous):

yea i know and thats how i solved for c2=12

OpenStudy (lalaly):

wait let me do it

OpenStudy (lalaly):

\[y(0)=2\]\[c_1+c_3=2\] \[y'(0)=5\]\[-0.5c_1+c_2=5\] \[y''(0)=6\]\[0.25c_1=6\]\[c_1=24\] \[24+c_3=2\]\[c_3=-22\]\[-0.5(24)+c_2=5\]\[c_2=17\]

OpenStudy (anonymous):

I got c2=-7 and thats where i went wrong! thanks soooo much!

OpenStudy (lalaly):

:D:D its ok, its confusing i know lol, ur welcome

OpenStudy (anonymous):

i frogot .5 was negative

OpenStudy (anonymous):

Thanks again! :)

OpenStudy (lalaly):

anytime:)

OpenStudy (anonymous):

can I ask another question?

OpenStudy (lalaly):

sure:D

OpenStudy (anonymous):

8y''+4y'+4y=0, where y(0)=4 and y'(0)=1

OpenStudy (lalaly):

u want the answers or the whole solution?

OpenStudy (anonymous):

the solution please

OpenStudy (anonymous):

but if you cant you can give the answer and I can try to work backwards

OpenStudy (lalaly):

ok give me 5 min let me work it out

OpenStudy (anonymous):

thanks sooooo much

OpenStudy (lalaly):

\[y=c_1e^{-\frac{x}{4}}\cos(\frac{\sqrt 7}{4}x)+c_2e^{-\frac{x}{4}}\sin(\frac{\sqrt 7}{4}x)\] i got \[c_1=4\]\[c_2=3.02\]

OpenStudy (anonymous):

ok let me check thanks again

OpenStudy (lalaly):

:):)

OpenStudy (anonymous):

woohoooo its correct! thanks sooo much

OpenStudy (lalaly):

hehe yaaay :D

OpenStudy (anonymous):

ok i have one last question but if you dont want to do it its totally fine! i had a bunch of roblems to do and these were the three i couldnt get so if its ok can i ask you another one? i feel bad for asking!

OpenStudy (lalaly):

lol dont feel bad, id love to help, why else would i be on openstudy

OpenStudy (anonymous):

true lol thanks soo much! here is the question:

OpenStudy (anonymous):

Find a linear homogeneous constant-coefficient differential equation with the general solution: y(x)=c1e^(2x) + c2Cos(9x) +c3Sin(9x) that has the form y'''+_________y''+_____y'+____y=0 the underscores are supposed to represent blanks lol

OpenStudy (lalaly):

i got off here by mistake :S let me do it again

OpenStudy (anonymous):

okie dokes! :)

OpenStudy (lalaly):

the solutions are 2,+/-9i

OpenStudy (lalaly):

so (r-2)(r^2+81)=0

OpenStudy (lalaly):

r^3-2r^2+81r-162=0

OpenStudy (lalaly):

y'''-2y''+81y'-162y=0

OpenStudy (anonymous):

how did you get the solutions?

OpenStudy (lalaly):

do u know how u find the general solution like when u get r=2,3 the general solution is y=c1e^2x+c_2e^3x and when u get complex solutions a+bi and a-bi the general solution is y=e^ax(c1cos(bx)+c2sin(bx)... u just do it back wards

OpenStudy (anonymous):

oh ok I got it! thnkssss sooo much!!! i really appreciate all yr help!! I also noticed that you joioned 3arabi open stidy group! r u taking arabic classes?

OpenStudy (anonymous):

study*

OpenStudy (lalaly):

lol no im half arab :D:D i speak arabic

OpenStudy (anonymous):

oh lol I was going to ask if you need any help because I also speak arabic and it's my minor

OpenStudy (lalaly):

oh wow really :D:D where are u from hehe

OpenStudy (anonymous):

I'm palestinian! hbu?

OpenStudy (lalaly):

o ana kmaaaaaaaaaaaaaan :D

OpenStudy (anonymous):

but if you're referring to city, I'm from Dearborn lol

OpenStudy (anonymous):

omg nooo waaayyy!!! how cool!

OpenStudy (lalaly):

hehe im from hebron hehe

OpenStudy (anonymous):

oh by dearborn i meant city in the US haha I'm from beit safafa in palestine

OpenStudy (anonymous):

do you live in palestine?

OpenStudy (lalaly):

oh lol i dont know abt US lol, ana 3ayshe bi newzealand

OpenStudy (anonymous):

ooooo keef new zealns? keef al; taqs hunak?

OpenStudy (anonymous):

zealand*

OpenStudy (lalaly):

hehe its summer so yaaaaaaaaay !! :D newzealand btjanen bmoot feha hehe, 3njad sodfe helwee:D:D iza bedik more help ask me :D

OpenStudy (anonymous):

3njad inti bit mooti feeha? lol hon it's winter bes ana behab al pelletaa3! bhab al thalj! and shukran bes khalasit wajabatee! thanks soo much! i really appreciate your help!

OpenStudy (anonymous):

i was supposed to say pelletaa3 not pelletaa3 lol i have no idea hot i typed that instead! lol sorry for the confusion

OpenStudy (anonymous):

pelletaaaa3 as in winter* i dont know why it keeps sending it as something else

OpenStudy (lalaly):

if u catch me on here at anytime ill be glad to help u o sara7a u made my day hehe i was a lil upset, bas u made me smile :D hehe Fhemt-ha

OpenStudy (anonymous):

thanks so much! and same to you! if you have any questions I'll be more than glad to help to the best of my ability! and I'm glad to have made your day! a smile is always better than a frown! :)

OpenStudy (lalaly):

hehe nice to meet you, esme lana btw

OpenStudy (anonymous):

nice to meet you lana! ismee Ranya

OpenStudy (lalaly):

oooo kman my favorite name ya salam hehe

OpenStudy (anonymous):

haha shukran!

OpenStudy (anonymous):

btasif bes lazim arooh leana lazam arooh ijeeb igrath min al mahalat! bye! and talk to you soon inshallah! :)Have a great day!! :)

OpenStudy (lalaly):

:D:D okay take care,, bye byee

OpenStudy (anonymous):

bye :)